# One Model to Rule them All: Towards Universal Segmentation for Medical Images with Text Prompts

### v2

@@ 5 Dataset
5.1 Domain Knowledge
To acquire textual knowledge, we mainly exploit two sources of domain knowledge: the Unified Medical Language System (UMLS) [^1] https://www.nlm.nih.gov/research/umls, a comprehensive medical knowledge graph consisting of concept definitions and their interrelations; and search engines, which are prompted to organize knowledge into a graph of the same format, specifically refined for the human anatomy corpus. Regarding visual knowledge, we have compiled 72 medical segmentation datasets, creating an atlas that covers over 497 anatomical structures of the human body. All sources of knowledge are systematically categorized, encompassing anatomies and concepts from all regions of the human body. Two examples of these sources are illustrated in Fig. 11 (a) and (b).
@@ 5.1 Domain Knowledge
Search engine generared knowledge. Although UMLS is widely acknowledged and adopted as a general medical knowledge corpus [109, 116, 54, 121], it lacks a fine-grained description of anatomy concepts critical for segmentation tasks. For example, for ‘liver’, the definition is ‘A large lobed glandular organ in the abdomen of vertebrates that is responsible for detoxification, metabolism, synthesis, and storage of various substances.’, which erases the morphological features and focuses on functionality. Meanwhile, more comprehensive knowledge on human anatomy may be scattered across various authoritative websites online, e.g., Wikipedia [^3] https://www.wikipedia.org, ScienceDirect [^4] https://www.sciencedirect.com/browse/journals-and-books?subject=medicine-and-dentistry, etc.. To harvest such knowledge, we select 6,502 anatomy concepts, and prompt an search engine  [^5] https://you.com to retrieve and summarize definitions for them. We use the following prompt template: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] Definition of xxx. Include the location, shape, appearance, structure and spatial relations to other anatomy structures. No need to include functionality. End with ‘END’. For illustration, the search engine referred to authority websites including Columbiasurgery [^6] https://columbiasurgery.org/liver/liver-and-its-functions, Hopkins Medicine [^7] https://www.hopkinsmedicine.org/health/conditions-and-diseases/liver-anatomy-and-functions and summerized definition for ‘liver’ as: ‘A large organ found in the upper right quadrant of the abdomen, it stands as the largest gland within the human body, with a weight of about 1.5 kilograms. This structure exhibits a reddish-brown hue and is cone or wedge-shaped … …’. While constructing the knowledge graph, we also adopt GPT4 [2] to extract 38,344 relations between anatomical structures in the generated information-dense definitions with the following prompt: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] This is the description of xxx. Please help me find its relations with other anatomical structures in radiological images. Summarize them with the template: Relation: xxx (relational preposition), Anatomical structure: xxx (name of another anatomical structure).
Visual Atlas. Segmentation datasets naturally provide visual features for anatomy concepts corresponding or complementary to the textual description, such as the texture, spatial location, shape and so on. Details on our collected segmentation datasets are described in Sec. 5.2. Here, we use them as a large-scale and diverse visual atlas library, and link the visual regions to corresponding concepts in textual knowledge graph, bridging the knowledge between visual and language modality.
5.2 Segmentation Dataset
To train our segmentation model with the ability to handle segmentation tasks of different targets, across various modalities and anatomical regions, we collect and integrate 72 diverse publicly available medical segmentation datasets, totaling 22,186 scans including both CT and MRI and 302,033 segmentation annotations spanning 8 different regions of the human body: Brain, Head and Neck, Upper Limb, Thorax, Abdomen, Plevis, and Lower Limb. The dataset is termed as SAT-DS. More details are present in Table LABEL:tab:dataset_details and Table LABEL:tab:dataset_links. Note that, some public datasets are not mutually exclusive, e.g., KITS23 and KITS21 [35], we thus only collect the latest version, to avoid redundancy and potential data leakage in train-test split.
@@ 5.2 Segmentation Dataset
6 Method
Towards building up our universal segmentation model by text prompt, we consider two main stages, namely, multimodal knowledge injection and universal segmentation training. In the following, we start by defining the problem scenario (Sec. 6.1). Then, we show how to structure multimodal medical knowledge and inject it into a text encoder (Sec. 6.2). Lastly, we employ the text encoder to guide our universal segmentation model training on SAT-DS dataset (Sec. 6.3).
6.1 Problem Formulation
@@
6.2 Multimodal Knowledge Injection
@@
  • Visual medical concept pair. To align with the segmentation task, we gather pairs consisting of a concept (can be either an anatomical structure or lesion) and its image atlas. Note that, multiple pairs could be extracted from a single image. These pairs share a similar format to the segmentation data, denoted as $(x_{i},y_{i};t_{i})$, where $x_{i}$ and $t_{i}$ are consistent with their definition in Sec. 6.1 and $y_{i}\in\mathbb{R}^{H\times W\times D\times 1}$ is a binary segmentation mask for $t_{i}$.
@@
6.3 Universal Segmentation Training
@@
Visual Encoding. We adopt the pre-trained visual encoder as explained in Sec. 6.2, Equa. 3, and continue training it.
@@ 7 Implement Details
Progressive Multimodal Knowledge Injection. For the stability of training, we implement the multimodal knowledge injection procedure introduced in Sec. 6.2 progressively. At the beginning, the text encoder is pretrained on the text knowledge via contrastive learning, i.e., the textual medical concept pairs. The maximal sequence length after tokenization is 256, for even longer text input, we apply random truncation to fully exploit the knowledge in the long text. After convergence, to align text representations and visual features, we apply contrastive learning between the finetuned text encoder and the visual encoder on the visual medical concept pair.
@@
8 Experiment Settings
8.1 Evaluation Datasets
@@
  • SAT-DS. As describe in Sec. 5.2, this contains all the 72 datasets, 497 classes from all human body regions, 22,186 image scans and 302,033 segmentation annotations.
@@
The detailed composition of SAT-DS and SAT-DS-Nano can be found in Table LABEL:tab:dataset_details. As there is no existing benchmark for evaluating the universal segmentation model, we randomly split each dataset into 80% for training and 20% for testing: (i) datasets may share the same images but with different classes. For example, Couinaud Liver provides fine-grained liver segmentation on a subset of MSD Hepatic Vessel. We carefully split the Couinaud Liver to make sure the test set will not be leaked in the train set of MSD Hepatic Vessel; (ii) scans of the same patient but different modalities are treated as different samples during training and evaluation. For example, MSD Prostate contains T2 and ADC scans of each patient. However, they share the same structure on the image. To avoid potential data leaking, we carefully split such datasets by patient id. Note that when involving segmentation datasets in the visual-language pretraining, we only use the training data to avoid potential data leaking. For datasets involved in SAT-DS-Nano, we keep their splits the same as in SAT-DS. The download link for each datasets can be found in Table LABEL:tab:dataset_links, and we will release our dataset splits to the research community for model comparison.
8.2 Baseline
We take nnU-Net [38] as a strong baseline for comparison, which specializes in a series of configurations for each dataset. For a comprehensive evaluation, we train one nnU-Net model on each of the datasets. Note that, following [106], we split Totalsegmentator into 6 subsets and train 6 different nnU-Net models. Similarly, datasets such as CHAOS with both MRI and CT images are treated as two different datasets. When training nnU-Nets on each dataset, we adopt a multi-class segmentation setting and deliver the masks of all categories in this dataset at once. We derive the optimal network architecture and pre-processing pipeline with the default setting of nnU-Net. The detailed network design is shown in Table LABEL:tab:nnunet_config. In summary, we train an assemble of 72 nnU-Nets, that are customized on each dataset and serve as the competitive baseline for evaluation. We adopt the latest official implementation of nnU-Net v2888https://github.com/MIC-DKFZ/nnUNet/ in practice.
8.3 Evaluation Protocols
Given our goal is to develop a universal medical segmentation model, this provides opportunities to evaluate novel perspectives in addition to the traditional evaluation per dataset. Specifically, we conduct the evaluations from three dimensions:
@@ 8.3 Evaluation Protocols
  • Dataset-wise Evaluation. Results of the classes within the same dataset are averaged to indicate the performance on this dataset. This is the same as the conventional evaluation protocol of specialist segmentation models trained on a single dataset, however, we believe this is less valuable for evaluating universal segmentation models.
8.4 Evaluation Metrics
@@ 9 Conclusion
In this paper, we promote the progress of medical universal segmentation with text as prompt and knowledge enhancement. We build up the largest and most comprehensive 3D medical segmentation datase, and the first multi-modal knowledge tree for human anatomy. In contrast to existing segmentation models that are limited in 2D input and box/point prompts, our model, SAT, enables text-driven universal segmentation on various radiologic modalities across 8 main human body parts, greatly reducing the usage difficulty of medical universal segmentation model in practice. The final solution we proposed, i.e., SAT-Pro, which only contains 447M paramters, while demonstrating comparable performance to 72 specialist nnU-Nets, with around 2.2B parameters trained individually on 72 datasets for segmenting 497 categories. We demonstrate its strong generalization capability to clinic data, as an off-the-shelf automatic segmentation method serving the doctor in analysis and diagnosis. Importantly, we show the potential of SAT as a powerful agent, paving the way to equip large language models with grounding capability.

### v3

@@ 4.1 Dataset Construction
4.1.1 Domain Knowledge
To acquire textual knowledge, we mainly exploit two sources of domain knowledge: the Unified Medical Language System (UMLS) [^1] https://www.nlm.nih.gov/research/umls, a comprehensive medical knowledge graph consisting of concept definitions and their interrelations; search engines, which are prompted to organize knowledge into a graph of the same format, specifically refined for the human anatomy corpus. Regarding visual knowledge, we have compiled 72 medical segmentation datasets, creating an atlas that covers over 497 anatomical structures of the human body. All sources of knowledge are systematically categorized, encompassing anatomies and concepts from all regions of the human body. Two examples of these sources are illustrated in Fig. 10 (a) and (b).
Domain Knowledge from Search Engine. Although UMLS is widely acknowledged and adopted as a general medical knowledge corpus [91, 96, 42, 100], it lacks a fine-grained description of anatomy concepts critical for segmentation tasks. For example, for ‘liver’, the definition is ‘A large lobed glandular organ in the abdomen of vertebrates that is responsible for detoxification, metabolism, synthesis, and storage of various substances.’, which erases the morphological features and focuses on functionality. Meanwhile, more comprehensive knowledge on human anatomy may be scattered across various authoritative websites online, e.g., Wikipedia [^3] https://www.wikipedia.org, ScienceDirect [^4] https://www.sciencedirect.com/browse/journals-and-books?subject=medicine-and-dentistry, etc.. To harvest such knowledge, we select 6,502 anatomy concepts, and prompt an search engine  [^5] https://you.com to retrieve and summarize definitions for them. We use the following prompt template: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] Definition of xxx. Include the location, shape, appearance, structure and spatial relations to other anatomy structures. No need to include functionality. End with ‘END’. For illustration, the search engine referred to authority websites including Columbiasurgery [^6] https://columbiasurgery.org/liver/liver-and-its-functions, Hopkins Medicine [^7] https://www.hopkinsmedicine.org/health/conditions-and-diseases/liver-anatomy-and-functions and summerized definition for ‘liver’ as: ‘A large organ found in the upper right quadrant of the abdomen, it stands as the largest gland within the human body, with a weight of about 1.5 kilograms. This structure exhibits a reddish-brown hue and is cone or wedge-shaped … …’. While constructing the knowledge graph, we also adopt GPT4 [2] to extract 38,344 relations between anatomical structures in the generated information-dense definitions with the following prompt: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] This is the description of xxx. Please help me find its relations with other anatomical structures in radiological images. Summarize them with the template: Relation: xxx (relational preposition), Anatomical structure: xxx (name of another anatomical structure).
Visual Atlas. Segmentation datasets naturally provide visual features for anatomy concepts corresponding or complementary to the textual description, such as the texture, spatial location, shape and so on. Details on our collected segmentation datasets are described in Sec. 4.1.2. Here, we use them as a large-scale and diverse visual atlas library, and link the visual regions to corresponding concepts in textual knowledge graph, bridging the knowledge between visual and language modality.
@@ 4.1.1 Domain Knowledge
4.1.2 Segmentation Dataset
To train our segmentation model with the ability to handle segmentation tasks of different targets, across various modalities and anatomical regions, we collect and integrate 72 diverse publicly available medical segmentation datasets, totaling 22,186 scans including both CT and MRI and 302,033 segmentation annotations spanning 8 different regions of the human body: Brain, Head and Neck, Upper Limb, Thorax, Abdomen, Plevis, and Lower Limb. The dataset is termed as SAT-DS. More details are present in Table 26 and Table 4. Note that, some public datasets are not mutually exclusive, e.g., KITS23 and KITS21 [24], we thus only collect the latest version, to avoid redundancy and potential data leakage in train-test split.
@@ 4.1.2 Segmentation Dataset
4.2 Universal Segmentation Prompted by Text
Towards building up our universal segmentation model by text prompt, we consider two main stages, namely, multimodal knowledge injection and universal segmentation training. In the following, we start by defining the problem scenario (Sec. 4.2.1). Then, we show how to structure multimodal medical knowledge and inject it into a text encoder (Sec. 4.2.2). Lastly, we employ the text encoder to guide our universal segmentation model training on SAT-DS dataset (Sec. 4.2.3).
4.2.1 Problem Formulation
@@
4.2.2 Multimodal Knowledge Injection
@@
  • Visual Medical Concept Pair. To align with the segmentation task, we gather pairs consisting of a concept (can be either an anatomical structure or lesion) and its image atlas. Note that, multiple pairs could be extracted from a single image. These pairs share a similar format to the segmentation data, denoted as $(x_{i},y_{i};t_{i})$, where $x_{i}$ and $t_{i}$ are consistent with their definition in Sec. 4.2.1 and $y_{i}\in\mathbb{R}^{H\times W\times D\times 1}$ is a binary segmentation mask for $t_{i}$.
@@
4.2.3 Universal Segmentation Training
@@
Visual Encoding. We adopt the pre-trained visual encoder as explained in Sec. 4.2.2, Equa. 3, and continue training it.
@@ 4.3 Implement Details
Progressive Multimodal Knowledge Injection. For the stability of training, we implement the multimodal knowledge injection procedure introduced in Sec. 4.2.2 progressively. At the beginning, the text encoder is pretrained on the text knowledge via contrastive learning, i.e., the textual medical concept pairs. The maximal sequence length after tokenization is 256, for even longer text input, we apply random truncation to fully exploit the knowledge in the long text. After convergence, to align text representations and visual features, we apply contrastive learning between the finetuned text encoder and the visual encoder on the visual medical concept pair.
@@
4.4 Evaluation Settings
4.4.1 Evaluation Datasets
@@
  • SAT-DS. As describe in Sec. 4.1.2, this contains all the 72 datasets, 497 classes from all human body regions, 22,186 image scans and 302,033 segmentation annotations.
@@
The detailed composition of SAT-DS and SAT-DS-Nano can be found in supplementary material Table 26 and Table 27. As there is no existing benchmark for evaluating the universal segmentation model, we randomly split each dataset into 80% for training and 20% for testing: (i) datasets may share the same images but with different classes. For example, Couinaud Liver provides fine-grained liver segmentation on a subset of MSD Hepatic Vessel. We carefully split the Couinaud Liver to make sure the test set will not be leaked in the train set of MSD Hepatic Vessel; (ii) scans of the same patient but different modalities are treated as different samples during training and evaluation. For example, MSD Prostate contains T2 and ADC scans of each patient. However, they share the same structure on the image. To avoid potential data leaking, we carefully split such datasets by patient id. Note that when involving segmentation datasets in the visual-language pretraining, we only use the training data to avoid potential data leaking. For datasets involved in SAT-DS-Nano, we keep their splits the same as in SAT-DS. The download link for each datasets can be found in Sec. 7, and we have released our dataset processing code and train-test splits to the research community for reproduction and benchmarking.
4.4.2 Baseline
Specialist Model. We take nnU-Net [27] as a strong baseline for comparison, which specializes in a series of configurations for each dataset. For a comprehensive evaluation, we train one nnU-Net model on each of the datasets. Note that, following [90], we split Totalsegmentator into 6 subsets and train 6 different nnU-Net models. Similarly, datasets such as CHAOS with both MRI and CT images are treated as two different datasets. When training nnU-Nets on each dataset, we adopt a multi-class segmentation setting and deliver the masks of all categories in this dataset at once. We derive the optimal network architecture and pre-processing pipeline with the default setting of nnU-Net. The detailed network design is shown in supplementary material Table 20 and Table 21. In summary, we train an assemble of 72 nnU-Nets, that are customized on each dataset and serve as the competitive baseline for evaluation. We adopt the latest official implementation of nnU-Net v2888https://github.com/MIC-DKFZ/nnUNet/ in practice. In zero-shot evaluation,
Interactive Segmentation Model. We take MedSAM [54] as a representative interactive universal segmentation model as baseline. MedSAM finetunes SAM [34] on 86 segmentation datasets, and supports 2D medical image segmentation with box prompts. We follow the official implement [^9] https://github.com/bowang-lab/MedSAM to process and infer image slice by slice, and calculate the metrics on the finally stacked 3D prediction. For each single target on a slice, to simulate box prompts towards it, we both take the minimum rectangle containing the ground truth segmentation (Tight), and following the official data augmentation procedure, randomly shift each corner up to 8% of the whole image resolution (Loose). In addition, we consider directly use the tight box prompts as prediction (Oracle Box).
4.4.3 Evaluation Protocols
Given our goal is to develop a universal medical segmentation model, this provides opportunities to evaluate novel perspectives in addition to the traditional evaluation per dataset. Specifically, we conduct the internal evaluations from three dimensions:
@@ 4.4.3 Evaluation Protocols
  • Dataset-wise Evaluation. Results of the classes within the same dataset are averaged to indicate the performance on this dataset. This is the same as the conventional evaluation protocol of specialist segmentation models trained on a single dataset.
4.4.4 Evaluation Metrics
@@ 5 Conclusion
In this paper, we promote the progress of medical universal segmentation with text as prompt and knowledge enhancement. We build up the largest and most comprehensive 3D medical segmentation dataset, and the first multi-modal knowledge tree for human anatomy. In contrast to existing segmentation methods that requires training and deploying a series of specialized models, or limited to box/point prompts, our model, SAT, enables text-driven universal segmentation on 3 radiological modalities, 497 anatomical structures or lesions, across 8 main human body parts in just one model, demonstrates great efficiency and flexibility in application. We proposed SAT-Nano (110M parameters) and SAT-Pro (447M parameters), demonstrating comparable performance to 72 specialist nnU-Nets, with around 2.2B parameters trained individually on each datasets, and showing even superior generalization ability on external datasets. Compared with the state-of-the-art interactive segmentation model, i.e., MedSAM, SAT achieves superior performance, scalability and robustness with text as prompts. SAT-Pro can also act as a foundation segmentation model, adapted to specific tasks for improved performance. We further demonstrate the usage of SAT as a powerful out-of-the-box agent for large language models, seamlessly providing grounding ability in clinical applications such as report generation. In summary, we believe our work can enhance versatile, efficient and easy-to-use medical segmentation in clinical applications, and contribute substantially to the development of generalist medcial artificial intelligence.
6 Code Availability
The code is available at https://github.com/zhaoziheng/SAT.
7 Data Availability of SAT-DS
The access to each dataset can be found in Tab. 4 and Tab. 5. The data process code to build SAT-DS and our train-test splits for reproducibility and benchmarking are available at https://github.com/zhaoziheng/SAT-DS.
Table 4: Download links of the 72 datasets in SAT-DS.
Dataset Download Link
AbdomenCT1K [57] https://github.com/JunMa11/AbdomenCT-1K
ACDC [9] https://humanheart-project.creatis.insa-lyon.fr/database/
AMOS CT [29] https://zenodo.org/records/7262581
AMOS MRI [29] https://zenodo.org/records/7262581
ATLASR2 [46] http://fcon_1000.projects.nitrc.org/indi/retro/atlas.html
ATLAS [72] https://atlas-challenge.u-bourgogne.fr
autoPET [16] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=93258287
Brain Atlas [81] http://brain-development.org/
BrainPTM [6] https://brainptm-2021.grand-challenge.org/
BraTS2023 GLI [60] https://www.synapse.org/#!Synapse:syn51514105
BraTS2023 MEN [36] https://www.synapse.org/#!Synapse:syn51514106
BraTS2023 MET [62] https://www.synapse.org/#!Synapse:syn51514107
BraTS2023 PED [32] https://www.synapse.org/#!Synapse:syn51514108
BraTS2023 SSA [3] https://www.synapse.org/#!Synapse:syn51514109
BTCV Abdomen [38] https://www.synapse.org/#!Synapse:syn3193805/wiki/217789
BTCV Cervix [39] https://www.synapse.org/#!Synapse:syn3193805/wiki/217790
CHAOS CT [31] https://chaos.grand-challenge.org/
CHAOS MRI [31] https://chaos.grand-challenge.org/
CMRxMotion [88] https://www.synapse.org/#!Synapse:syn28503327/files/
Couinaud [83] https://github.com/GLCUnet/dataset
COVID-19 CT Seg [55] https://github.com/JunMa11/COVID-19-CT-Seg-Benchmark
CrossMoDA2021 [14] https://crossmoda.grand-challenge.org/Data/
CT-ORG [75] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=61080890
CTPelvic1K [48] https://zenodo.org/record/4588403#ẎEyLq_0zaCo
DAP Atlas [28] https://github.com/alexanderjaus/AtlasDataset
FeTA2022 [67] https://feta.grand-challenge.org/data-download/
FLARE22 [56] https://flare22.grand-challenge.org/
FUMPE [59] https://www.kaggle.com/datasets/andrewmvd/pulmonary-embolism-in-ct-images
HAN Seg [69] https://zenodo.org/record/
HECKTOR2022 [4] https://hecktor.grand-challenge.org/Data/
INSTANCE [45] https://instance.grand-challenge.org/Dataset/
ISLES2022 [25] http://www.isles-challenge.org/
KiPA22 [22] https://kipa22.grand-challenge.org/dataset/
KiTS23 [1] https://github.com/neheller/kits23
LAScarQS2022 Task 1 [43] https://zmiclab.github.io/projects/lascarqs22/data.html
LAScarQS2022 Task 2 [43] https://zmiclab.github.io/projects/lascarqs22/data.html
LNDb [68] https://zenodo.org/record/7153205#Ẏz_oVHbMJPZ
Table 5: (Continued) Download links of the 72 datasets in SAT-DS.
Dataset Download Link
LUNA16 [82] https://luna16.grand-challenge.org/
MM-WHS CT [103] https://mega.nz/folder/UNMF2YYI#1cqJVzo4p_wESv9P_pc8uA
MM-WHS MR [103] https://mega.nz/folder/UNMF2YYI#1cqJVzo4p_wESv9P_pc8uA
MRSpineSeg [66] https://www.cg.informatik.uni-siegen.de/en/spine-segmentation-and-analysis
MSD Cardiac [5] http://medicaldecathlon.com/
MSD Colon [5] http://medicaldecathlon.com/
MSD HepaticVessel [5] http://medicaldecathlon.com/
MSD Hippocampus [5] http://medicaldecathlon.com/
MSD Liver [5] http://medicaldecathlon.com/
MSD Lung [5] http://medicaldecathlon.com/
MSD Pancreas [5] http://medicaldecathlon.com/
MSD Prostate [5] http://medicaldecathlon.com/
MSD Spleen [5] http://medicaldecathlon.com/
MyoPS2020 [70] https://mega.nz/folder/BRdnDISQ#FnCg9ykPlTWYe5hrRZxi-w
NSCLC [8] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=68551327
Pancreas CT [78] https://wiki.cancerimagingarchive.net/display/public/pancreas-ct
Parse2022 [50] https://parse2022.grand-challenge.org/Dataset/
PDDCA [74] https://www.imagenglab.com/newsite/pddca/
PROMISE12 [47] https://promise12.grand-challenge.org/Details/
SEGA [73] https://multicenteraorta.grand-challenge.org/data/
SegRap2023 Task1 [51] https://segrap2023.grand-challenge.org/
SegRap2023 Task2 [51] https://segrap2023.grand-challenge.org/
SegTHOR [37] https://competitions.codalab.org/competitions/21145#learn_the_details
SKI10 [41] https://ambellan.de/sharing/QjrntLwah
SLIVER07 [23] https://sliver07.grand-challenge.org/
ToothFairy [13] https://ditto.ing.unimore.it/toothfairy/
TotalSegmentator Cardiac [90] https://zenodo.org/record/6802614
TotalSegmentator Muscles [90] https://zenodo.org/record/6802614
TotalSegmentator Organs [90] https://zenodo.org/record/6802614
TotalSegmentator Ribs [90] https://zenodo.org/record/6802614
TotalSegmentator Vertebrae [90] https://zenodo.org/record/6802614
TotalSegmentator V2 [90] https://zenodo.org/record/6802614
VerSe [80] https://github.com/anjany/verse
WMH [35] https://wmh.isi.uu.nl/
WORD [52] https://github.com/HiLab-git/WORD